Which savings account is better, one paying 6 % with yearly compounding or one paying 5.9 % with daily compounding?
Answer #1

EAR=(1+APR/m)^m-1
where m=compounding periods

1st case:

EAR=(1+0.06/1)^1-1

=6%

2nd case:

EAR=(1+0.059/365)^365-1[Taking 365 days in a year]

=6.08%(Approx)

Hence 5.9% with daily compounding is better.